OpenStudy (anonymous):

You get a discount of 20% on a 27.50 ticket.How much will your ticket cost? Desribe your method,and explain why you choose that answer.

OpenStudy (anonymous):

so we have a 27.50 ticket ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and we are getting a 20% discount. How would you write 20% in like mathematical terms?

OpenStudy (anonymous):

.20

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

So it is basically 20% off 27.50 Do u remember how we wld set up the equation?

OpenStudy (anonymous):

LOL now that I am reading it I am realizing that it is different than yesterday

OpenStudy (anonymous):

.20=n27.50

OpenStudy (anonymous):

uhhhh no

OpenStudy (anonymous):

It is different than yesterday

OpenStudy (anonymous):

ok so i dont know how you learnt it but I will show you both ways

OpenStudy (anonymous):

why just one way

OpenStudy (anonymous):

ok so basically we are asking what is 20% of 27.50 27.50*.20=5.5

OpenStudy (anonymous):

so we see that 20% of 27.50 is equal to 5.50

OpenStudy (anonymous):

did you follow?

OpenStudy (anonymous):

so we are saying 20% of the amount the ticket would cost id $5.50

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

=

OpenStudy (anonymous):

no I am just correcting what i wrote b4 :D

OpenStudy (anonymous):

ok so now our equation says 20% off of 27.50

OpenStudy (anonymous):

so meaning that we have to remove the 20% from 27.50

OpenStudy (anonymous):

so basically the 20% of 27.50 = 5.50

OpenStudy (anonymous):

and so we subtract 5.50 from 27.50 27.50-5.50= 22 so when you get 20% off the ticket the ticket will equal $22

OpenStudy (anonymous):

LOL i tried to be clear but i doubt it

OpenStudy (anonymous):

Star how old are you?

OpenStudy (anonymous):

u dont wanna know

OpenStudy (anonymous):

LOL ok but what didnt u understand

OpenStudy (anonymous):

Star?

OpenStudy (anonymous):

i understand the answer is 22 right?

OpenStudy (anonymous):

ya lol I hope you understand how I got that. There is a simpler method but I only can teach u it if you understand this method correctly and it makes sense in your brain. So work on a few of these problems and when you are comfortable with it I can teach u a method that is only one step

OpenStudy (anonymous):

Lets take another example. A t-shirt costs $50 dollars. There is a sale and everything in the store is 45% off? What would the price of the t-shirt be?

OpenStudy (anonymous):

55

OpenStudy (anonymous):

So basically we would like to know what 45% of $50 is

OpenStudy (anonymous):

so we would set up our equation like this .45*50=? What is the answer?

OpenStudy (anonymous):

2250

OpenStudy (anonymous):

,45*50=22.5 This means that 45% of $50 is $22.50 Now since it is 45% off we have to subtract that amount from the original amount 50-22.50=27.50 so The t-shirt would cost $27.50

OpenStudy (anonymous):

sorry i didnt see that star but ya 22.50 is the correct answer :D

OpenStudy (anonymous):

If there is a 20% discount, then you will have to pay 80% of the $27.50 ticket price. 0.8 * 27.50 = $22.00